(32%) Problem 2: The magnetic field of a proton is relatively close to that of a circular current loop 0.650 x 10 m in radius carrying 1.05 x 104 A. An MRI machine is required to be able to manipulate these fields. Go 50% Part (a) To see why an MRI utilizes iron to increase the magnetic field generated by a coil, calculate the current needed in a 400 loop-per-meter circular coil 0.660 m in radius to generate a 1.20 T field (typical of an MRI instrument at its center with no iron insert.
Answer #1

Magnetic field at the center of the loop is B = mu_o*i*N/(2*R) = 1.2

mu_o*i*N/(2*R) = 1.2

(4*3.142*10^-7*i*400)/(2*0.66) = 1.2

current i= 3.15*10^3 A